Solve the recurrence relation a_n-4a_n-1+5a_n-2-2a_n-3=1+2^n
1
Expert's answer
2020-05-24T15:55:11-0400

"a_n-4a_{n-1}+5a_{n-2}-2a_{n-3}=1+2^n"

"a_n=4a_{n-1}-5a_{n-2}+2a_{n-3}+(1+2^n)"


Characteristic equation: "r^3-4r^2+5r-2=(r-1)^2(r-2)=0, \ \ r_1=1, r_2=2"

Then "a^{(0)}_n=(c_1+c_2n)r_1^n+c_3r_2^n=c_1+c_2n+c_32^n" is a solution of "a_n=4a_{n-1}-5a_{n-2}+2a_{n-3}"


We need do find particular solution of "a_n=4a_{n-1}-5a_{n-2}+2a_{n-3}+F(n), \ \ F(n)=(1+2^n)"

1 and 2 are characteristic roots of multiplicity 2 and 1 respectively.

Then particular solution has form of "a^{(p)}_n=n^2(b_tn^t+b_{t-1}n_{t-1}+...+b_1n+b_0)1^n+ n(d_kn^k+d_{k-1}n_{k-1}+...+d_1n+d_0)2^n"

Suppose that "a^{(p)}_n=n^2 b_0\times 1^n+nd_02^n=b_0n^2+d_0n2^n" and substitute it:

"b_0n^2+d_0n2^n=4(b_0(n-1)^2+d_0(n-1)2^{n-1})-5(b_0(n-2)^2+d_0(n-2)2^{n-2})+2(b_0(n-3)^2+d_0(n-3)2^{n-3})+1+2^n=b_0(n^2+2)+d_0n2^n-d_02^{n-2}+1+2^n"


"b_0n^2+d_0n2^n=b_0(n^2+2)+d_0n2^n-d_02^{n-2}+1+2^n"

"(2b_0+1)+(2^n-d_02^{n-2})=0, \ \ \forall n\in\mathbb{N}"

"b_0=-1/2, \ \ d_0=4"

Solution of the recurrence relation is sum of "a^{(0)}_n" and "a^{(p)}_n" .

We have the following solution of the recurrence relation:

"a_n=c_1+c_2n+c_32^n-1\/2n^2+n2^{n+2}" .



Answer: "a_n=c_1+c_2n+c_32^n-1\/2n^2+n2^{n+2}."


The following source is used:

https://courses.ics.hawaii.edu/ReviewICS241/morea/counting/RecurrenceRelations2-QA.pdf